# To Trust or Not to Trust: On Calibration in ML-based Resource Allocation for Wireless Networks

### v1

@@ Abstract
Index Terms:
Custom loss function, Isotonic regression, Machine learning, ML calibration, Outage prediction, Outage loss function, Outage probability, Post-processing calibration, Platt scaling, Resource allocation, Reproducible AI, Trustworthy AI.
@@ I Introduction
How should we interpret a neural network’s (NN) confidence score? If a NN assigns a confidence of 0.4 to the question “Will this communication channel fail in the future?” what does that really mean? What if it instead returns 0.000001? These questions highlight the importance of calibration—the extent to which a network’s confidence estimates align with the true probability of an event occurring.
@@
Accurate probability estimates are essential for informed decisions, yet NNs often produce confidence scores that do not accurately reflect true probabilities, a phenomenon known as poor calibration. When a model is miscalibrated, it may systematically overestimate or underestimate the likelihood of an event, leading to unreliable decision-making. Proper calibration aligns predicted probabilities with actual outcomes, enhancing both the reliability and usability of machine learning (ML)-driven predictions [1].
@@ I-A2 Custom Loss Functions for ML-Driven Wireless Systems
Most ML-based communication models highlighted previously, rely on traditional loss functions, such as binary cross-entropy (BCE) and mean squared error (MSE), to optimize model performance. While these loss functions are effective for general classification and regression tasks, they do not always align with the specific requirements of wireless networks [41]. In particular, conventional training objectives often fail to capture communication-specific constraints, such as OPs, resource efficiency, and latency guarantees. This misalignment can lead to suboptimal performance, prompting research into more specialized optimization approaches.
@@ I-A3 ML Calibration
Regardless of the loss function used, calibration remains a key challenge in ML-driven communication systems. To address this issue, Bayesian learning [51, 52] models epistemic uncertainty through parameter distributions to improve probabilistic predictions. It has been applied to a range of communication tasks, including network resource allocation [53, 54, 55], massive MIMO detection [56, 57, 58], and channel estimation [59]. While exact Bayesian learning offers formal calibration guarantees, these rely on strong assumptions. Specifically, they assume that the model is well specified, meaning the NN has enough capacity to represent the ground-truth data generation process, and that its predictive distribution for continuous outputs is unimodal. These assumptions may not be satisfied in cases where sufficient data is not available. In [60], the authors explored the application of robust Bayesian learning to wireless communication systems, focusing on applications such as automated modulation classification, RSSI-based localization, as well as channel modeling and simulation. However, this approach does not offer formal calibration guarantees. Conformal prediction [61] is a general framework for constructing set predictors that offer formal, distribution-free calibration guarantees and has been applied to problems such as symbol demodulation and modulation classification [62]. More recently, it has been explored as a tool to support uncertainty quantification through pre-deployment calibration, online monitoring, and post-deployment diagnostic analysis [63]. Another widely used approach is post-processing calibration, which adjusts confidence scores using a validation set. Common techniques include Platt scaling, isotonic regression, beta calibration for binary classification, and temperature or Dirichlet scaling for multiclass settings [64, 65, 66, 67, 68, 69, 70].
@@ I-B Contributions
    We demonstrate that in the infinite resource limit, the OP of the system, when the perfectly calibrated model is used equals, its expected confidence, conditioned on it being less than its classification threshold.
@@
    We show that when only one resource is available, and the model is perfectly calibrated, the model’s expected output is equal to the resource’s underlying OP.
@@
    Our theoretical analysis shows that, for well-calibrated models, there is a straightforward approach for choosing a classification threshold to achieve a desired system OP. For example, when a model is well calibrated and the number of available resources is sufficiently large, system designers need only choose a classification threshold equal to their desired OP. The resulting system OP will then be less than their selected threshold.
@@
    We show that well-calibrated models form a special case of a more general class of model that necessarily improves OP. Specifically, we establish a monotonicity condition that accuracy-confidence functions must satisfy in order to improve system OP.
@@
  6. 6.
    A rigorous simulation-based analysis in a Rayleigh fading environment using post-processing calibration techniques is conducted, demonstrating the theoretical findings above.
  7. 7.
    Finally, we present a preliminary proof of concept demonstrating the application of the resource allocation technique and the outage loss function used in this work and in [71]. This demonstration is intended for a broader audience.
This work builds on [2, 72] and presents a significantly more comprehensive study of the outage predictor calibration within this system. We analyze how calibration affects the classification threshold, enabling more effective resource selection, while also incorporating the additional contributions outlined above. For reproducibility purposes, we have made our code publicly available [^1] https://github.com/ML4Comms/greedy-resource-allocation-outage-classification [71]..
The remainder of this paper is organized as follows: Section II revisits the resource allocation strategy incorporating an outage predictor, as proposed in [46]. It then introduces the OLF used for its training, before providing an overview of calibration concepts. Section III discusses the key theoretical results of this work. Section IV begins with a description of data generation process, outlines the essential metrics for evaluating calibration performance, and provides an overview of the post-processing calibration techniques implemented in this paper. Section V presents the numerical results that demonstrate the findings established in Section III, and discusses the proof of concept for the resource allocation technique introduced in Section II. Finally, Section VI concludes the paper with some finishing remarks.
@@ II-A Channel Model for a Resource
Each resource $j\in\mathcal{R}$ has a fluctuating channel state $h_{j}(t)\in~{}\mathbb{C}$, with channel states assumed to be independent and identically distributed (i.i.d.) across distinct resources. Furthermore, the correlation between $h_{j}(t)$ and $h_{j}(t+l)$ decreases as $l\to\infty$. The user is equipped with an ML-based outage predictor that assists in selecting the appropriate resource by learning these temporal correlation patterns. For $k,l\in\mathbb{N}$, the vectors of past and future channel states at time $t$ are given by $\accentset{\leftarrow}{H}_{j}^{k}{(t)}\triangleq\left[h_{j}{(t-k+1)},\cdots,h_{j}{(t)}\right]^{T}$, and $\accentset{\rightarrow}{H}_{j}^{l}(t)\triangleq\left[h_{j}(t+1),\cdots,h_{j}(t+l)\right]^{T}$, respectively.
The capacity $C\left({\accentset{\rightarrow}{H}_{j}^{l}}{(t)}\right)\in\mathbb{R}^{+}$ measures resource $j$’s ability to support communication from time $t+\!1$ to $t+\!l$. For a quasi-static Gaussian channel it is given by [73, eq. (5.80)]:
@@
where $\mathbb{P}[\cdot]$, represents the ground-truth distribution of the observation state sequences $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$.
@@ II-B ML-based Resource Allocation
In order to quantify the calibration quality of the predictor, we adopt the standard characterization in terms of reliability diagrams. A reliability diagram is a visual representation of model calibration, plotting the observed accuracy as a function of confidence [1]. In this context, the accuracy-confidence function $A_{j}(q)$ describes the probability of an outage event given the predictor provides an outage confidence of $q$. For our outage predictor $Q\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$, the accuracy-confidence function is formally defined as
@@
Illustrative plots of the functions $A_{j}(q)$ and $F_{j}(q)$ are shown in Fig. 1. The top figure shows three accuracy-confidence curves corresponding to the extreme cases: perfect calibration $\left(A_{j}(q)=q\right)$, strong under-confidence $\left(A_{j}(q)>q\right)$, and strong over-confidence $\left(A_{j}(q)<q\right)$, for all $q\in[0,1]$. The bottom figure shows two examples of confidence distributions, where lower $F_{j}(q)$ indicates a model that tends to produce larger fraction of confident decisions.
@@
is learned solely from a held-out calibration dataset $\mathcal{D}$. Here, $\mathcal{C}\bigl{(}Q;\mathcal{D}\bigr{)}$ represents the calibration procedure, and $r\circ Q$ represents function composition i.e., the calibrated predictor is obtained by applying $r$ to the output of $Q$. Since $Q^{c}\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$ is obtained by (possibly stochastic) processing of $Q\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$, the data processing inequality implies
@@ II-C Greedy Resource Allocation
We adopt the resource allocation strategy from [46], in which the user scans the available resources $\mathcal{R}$ and selects a resource $j^{\prime}$ based on whether the outage predictor $Q^{\star}\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$ indicates successful communication. The predictor sequentially scans resources =$\mathcal{R}=\{1,2,\dots,|\mathcal{R}|\}$, stopping once it predicts successful communication for a given resource. If no resource meets this condition, the final resource is selected.
@@
~$\tilde{\mathcal{R}}({q_{th}})=\left\{j\in\mathcal{R}~{}\text{s.t.}~{}Q^{\star}\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right);\Theta)\leq q_{th}\right\},$ (10)
@@
Let $P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})\in\{P_{|\mathcal{R}|}(\gamma_{th},q_{th}),\;P^{c}_{|\mathcal{R}|}(\gamma_{th},q_{th})\}$ denote the system’s OP when using either the uncalibrated predictor $Q$ or the calibrated predictor $Q^{c}$, respectively. These are formally defined as:
@@
where $j^{\prime}(q_{th})$ is the selected resource index according to (11). Similarly, in the infinite-resource limit, as derived in [46], the OP is given by:
@@
Following [46, eq. (13)] and using the law of total probability, the OP in (12), expressed in terms of the predictor’s confidence distribution function and accuracy-confidence function is given by:
@@
$+\!\mathbb{E}\!\left[A_{j}\left(Q^{\star}\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)\right)\mid\!Q^{\star}\left(\!\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)\!\leq q_{th}\right]$
×$\times\left(1-(1-F_{j}(q_{th}))^{|\mathcal{R}|-1}\right).$ (14)
@@
The OLF, customized for this resource allocation system, is defined in [46, eq. (29)]. This loss function is utilized to train the outage predictor, enabling the generation of the results presented in Section V. For brevity, we drop the dependence of the sequences $\accentset{\leftarrow}{H}_{j}^{k}{(t)}$ and $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$ on $t$, as well as the dependence on $j$ for $P\triangleq P_{j}$, accuracy-confidence function $A(q)\triangleq A_{j}(q)$, and confidence distribution $F(q)\triangleq F_{j}(q)$. Additionally, where appropriate, we drop the dependence of the predictor’s confidence $Q^{\star}(\cdot)$ and of $C(\cdot)$ on $\accentset{\leftarrow}{H}_{j}^{k}{(t)}$ and $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$, respectively. We denote by $Q^{\star}$ a random variable with the same distribution as any confidence level $Q^{\star}$ assigned by the predictor to a resource $j$.
Before moving on to the next section, we reiterate our problem statement here. This paper builds on [2], which analyzed the general case where the outage predictor may or may not be well calibrated for an ML-assisted resource allocation system. In such cases, the predictor’s confidence scores may not accurately reflect the true probability of outages. Central to our study is the characterization of the system’s OP under perfect calibration which helps guide the choice of the classification threshold required to achieve a target OP when sufficient resources are available. We then apply post-processing calibration techniques to validate our theoretical findings.

### v3

@@ I Introduction
How should we interpret a neural network’s (NN) confidence score? If a NN assigns a confidence of 0.4 to the question “Will this communication channel fail in the future?” what does that really mean? What if it instead returns 0.000001? These questions highlight the importance of calibration—the extent to which a network’s confidence estimates align with the true probability of an event occurring. In essence, in machine learning (ML)-based systems, calibration concerns the process of adjusting a model’s predicted probabilities to correctly reproduce the true likelihood of an event.
@@
Accurate probability estimates are essential for informed decisions, yet NNs often produce confidence scores that do not accurately reflect true probabilities, a phenomenon known as poor calibration. When a model is miscalibrated, it may systematically overestimate or underestimate the likelihood of an event, leading to unreliable decision-making. Proper calibration aligns predicted probabilities with actual outcomes, enhancing both the reliability and usability of ML-driven predictions [3].
@@ I-A2 Custom Loss Functions for ML-Driven Wireless Systems
Most ML-based communication models highlighted previously rely on traditional loss functions, such as binary cross-entropy (BCE) and mean squared error (MSE), to optimize model performance. While these loss functions are effective for general classification and regression tasks, they do not always align with the specific requirements of wireless networks [43]. In particular, conventional training objectives often fail to capture communication-specific constraints, such as OPs, resource efficiency, and latency guarantees. This misalignment can lead to suboptimal performance, prompting research into more specialized optimization approaches.
@@ I-A3 ML Calibration
Regardless of the loss function used, calibration remains a key challenge in ML-driven communication systems. To address this issue, Bayesian learning [53, 54] models epistemic uncertainty through parameter distributions to improve probabilistic predictions. It has been applied to a range of communication tasks, including network resource allocation [55, 56, 57], massive MIMO detection [58, 59, 60], and channel estimation [61]. While exact Bayesian learning offers formal calibration guarantees, these rely on strong assumptions. Specifically, they assume that the model is well-specified, meaning the NN has enough capacity to represent the ground-truth data generation process, and that its predictive distribution for continuous outputs is unimodal. These assumptions may not be satisfied in cases where sufficient data is not available. In [62], the authors explored the application of robust Bayesian learning to wireless communication systems, focusing on applications such as automated modulation classification, RSSI-based localization, as well as channel modeling and simulation. However, this approach does not offer formal calibration guarantees. Conformal prediction [63] is a general framework for constructing set predictors that offer formal, distribution-free calibration guarantees and has been applied to problems such as symbol demodulation and modulation classification [64]. More recently, it has been explored as a tool to support uncertainty quantification through pre-deployment calibration, online monitoring, and post-deployment diagnostic analysis [65]. Another widely used approach is post-processing calibration, which adjusts confidence scores using a validation set. Common techniques include Platt scaling, isotonic regression, beta calibration for binary classification, and temperature or Dirichlet scaling for multiclass settings [66, 67, 68, 69, 70, 71, 72].
@@ I-B Contributions
    We show that for perfectly calibrated models, the system OP—in the infinite-resource limit—equals the model’s expected confidence, conditioned on it being less than its classification threshold, and equals the resource’s underlying OP when only a single resource is available.
@@
    We conduct a rigorous simulation-based analysis in a Rayleigh fading environment using post-processing calibration techniques, demonstrating our theoretical finding that well-calibrated models enable straightforward classification threshold selection. For instance, for a well-calibrated model with a sufficiently large number of resources, system designers need only choose a classification threshold equal to their desired OP, resulting in a system OP that is lower than the selected threshold.
@@
    We show that well-calibrated models form a special case of a more general class of models that necessarily improves OP. Specifically, we establish a monotonicity condition that accuracy-confidence functions must satisfy in order to improve system OP.
@@
    Finally, we present a preliminary proof of concept demonstrating the application of the resource allocation technique and the OLF used in this work, and in [73]. This demonstration is intended for a broader audience.
This work builds on [1] and presents a significantly more comprehensive study of the outage predictor calibration within this system. A small part of the simulation work established here, appears in [2]. In this paper, we analyze how calibration affects the classification threshold, enabling more effective resource selection, while also incorporating the additional contributions outlined above. For reproducibility purposes, we have made our code publicly available [^1] https://github.com/ML4Comms/greedy-resource-allocation-outage-classification [73]..
The remainder of this paper is organized as follows: Section II revisits the resource allocation strategy incorporating an outage predictor, as proposed in [48]. It then introduces the OLF used for its training, before providing an overview of calibration concepts. Section III discusses the key theoretical results of this work. Section IV begins with a description of the data generation process, outlines the essential metrics for evaluating calibration performance, and provides an overview of the post-processing calibration techniques implemented in this paper. Section V presents the numerical results that demonstrate the findings established in Section III, and discusses the proof of concept for the resource allocation technique introduced in Section II. Finally, Section VI concludes the paper with some finishing remarks.
@@ II-A Channel Model for a Resource
Each resource $j\in\mathcal{R}$ has a fluctuating channel state $h_{j}(t)\in\mathbb{C}$, with channel states assumed to be independent and identically distributed (i.i.d.) across distinct resources. Furthermore, the correlation between $h_{j}(t)$ and $h_{j}(t+l)$ decreases as $l\to\infty$. The user is equipped with an ML-based outage predictor that assists in selecting the appropriate resource by learning these temporal correlation patterns. For $k,l\in\mathbb{N}$, the vectors of past and future channel states at time $t$ are given by $\accentset{\leftarrow}{H}_{j}^{k}{(t)}\triangleq\left[h_{j}{(t-k+1)},\cdots,h_{j}{(t)}\right]^{T}$, and $\accentset{\rightarrow}{H}_{j}^{l}(t)\triangleq\left[h_{j}(t+1),\cdots,h_{j}(t+l)\right]^{T}$, respectively.
The capacity $C\left({\accentset{\rightarrow}{H}_{j}^{l}}{(t)}\right)\in\mathbb{R}^{+}$ measures resource $j$’s ability to support communication from time t+1t+\!1 to t+lt+\!l. For a quasi-static Gaussian channel, it is given by [74, eq. (5.80)]:
@@
where $\mathbb{P}[\cdot]$ represents the ground-truth distribution of the observation state sequences $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$.
@@ II-B ML-based Resource Allocation
In this study, calibration is defined as the ML predictor’s confidence matching the true outage frequency, with perfect calibration achieved when predictions with confidence $q$ correspond to an actual outage frequency of $q$. In order to quantify the calibration quality of the predictor, we adopt the standard characterization in terms of reliability diagrams. A reliability diagram is a visual representation of model calibration, plotting the observed accuracy as a function of confidence [3]. In this context, the accuracy-confidence function $A_{j}(q)$ describes the probability of an outage event given that the predictor provides an outage confidence of $q$. For our outage predictor $Q\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$, the accuracy-confidence function is formally defined as
@@
Illustrative plots of the functions $A_{j}(q)$ and $F_{j}(q)$ are shown in Fig. 1. The top figure shows three accuracy-confidence curves corresponding to the extreme cases: perfect calibration $\left(A_{j}(q)=q\right)$, strong under-confidence $\left(A_{j}(q)>q\right)$, and strong over-confidence $\left(A_{j}(q)<q\right)$, for all $q\in[0,1]$. The bottom figure shows two examples of confidence distributions, where a lower $F_{j}(q)$ indicates a model that tends to produce a larger fraction of confident decisions.
@@
is learned solely from a held-out calibration dataset $\mathcal{D}$. Here, $\mathcal{C}\bigl(Q;\mathcal{D}\bigr)$ represents the calibration procedure, and $r\circ Q$ represents function composition, i.e., the calibrated predictor is obtained by applying $r$ to the output of $Q$. Since $Q^{c}\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$ is obtained by (possibly stochastic) processing of $Q\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$, the data processing inequality implies
@@ II-C Greedy Resource Allocation
We adopt the resource allocation strategy from [48], in which the user scans the available resources $\mathcal{R}$ and selects a resource $j^{\prime}$ based on whether the outage predictor $Q^{\star}\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$ indicates successful communication. The predictor sequentially scans resources $\mathcal{R}=\{1,2,\dots,|\mathcal{R}|\}$, stopping once it predicts successful communication for a given resource. If no resource meets this condition, the final resource is selected.
@@
$\tilde{\mathcal{R}}({q_{th}})=\left\{j\in\mathcal{R}~\text{s.t.}~Q^{\star}\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)};\Theta\right)\leq q_{th}\right\},$ (10)
@@
Let $P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})\in\{P_{|\mathcal{R}|}(\gamma_{th},q_{th}),\;P^{c}_{|\mathcal{R}|}(\gamma_{th},q_{th})\}$ denote the OP of a system with $|\mathcal{R}|$ resources, when using either the uncalibrated predictor $Q$ or the calibrated predictor $Q^{c}$, respectively. These are formally defined as:
@@
where $j^{\prime}(q_{th})$ is the selected resource index according to (11). From [48, eq. (13)], the OP in (12) can be expressed as:
$P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})$ $={P_{j}(\gamma_{th})\left(1-F_{Q^{\star}}(q_{th})\right)^{|\mathcal{R}|-1}}$
$\thinspace+\!{P^{\star}_{\infty}(\gamma_{th},q_{th})\!\left(1-\!\left(1-\!F_{Q^{\star}}(q_{th})\right)^{|\mathcal{R}|-1}\!\right)},$ (13)
where $F_{Q^{\star}}(q_{th})$ is the cumulative distribution function of the predictor’s output $Q^{\star}$, and $P^{\star}_{\infty}(\gamma_{th},q_{th})$ is the OP of a system with an infinite number of resources, given by:
@@
Using (13), and the law of total probability, the OP can also be expressed in terms of the predictor’s confidence distribution function and the accuracy-confidence function. From (13), the term $(1-F_{Q^{\star}}(q_{th}))^{|\mathcal{R}|-1}$ represents the probability that none of the resources meets the condition $Q^{\star}\left(\!\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)\!\leq q_{th}$, in which case the final resource is selected, and the OP is given by $P_{j}(\gamma_{th})$. Conversely, the term $1-(1-F_{Q^{\star}}(q_{th}))^{|\mathcal{R}|-1}$ represents the probability that at least one of the resources satisfies the condition $Q^{\star}\left(\!\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)\!\leq q_{th}$. In this case, the OP is given by the expected accuracy $\mathbb{E}\left[A_{j}\left(Q^{\star}\left(\accentset{\leftarrow}{H}_{j}^{k}(t)\right)\right)\,\middle|\,Q^{\star}\left(\accentset{\leftarrow}{H}_{j}^{k}(t)\right)\leq q_{th}\right]$. Note that, due to the i.i.d. assumption on channel states across resources, the confidence distribution $F_{j}(q_{th})$ is identical for all $j\in\mathcal{R}$. Therefore, $F_{Q^{\star}}(q_{th})\equiv F_{j}(q_{th})$ for any resource $j$, and we use the notation $F_{j}(q_{th})$ in what follows. Thus, the OP is given by:
@@
$+\!\mathbb{E}\!\left[A_{j}\left(Q^{\star}\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)\right)\middle|Q^{\star}\left(\!\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)\!\leq q_{th}\right]$
$\times\left(1-(1-F_{j}(q_{th}))^{|\mathcal{R}|-1}\right).$ (15)
Equation (15) captures the relationship between the system’s OP, the accuracy-confidence function, and the confidence distribution function. This equation is further used in Theorem 3 (see Appendix A-C), which established a monotonicity condition that the accuracy-confidence function must satisfy in order to improve system OP.
@@
The OLF, customized for this resource allocation system, is defined in [48, eq. (29)]. This loss function is utilized to train the outage predictor, enabling the generation of the results presented in Section V. For brevity, we drop the dependence of the sequences $\accentset{\leftarrow}{H}_{j}^{k}{(t)}$ and $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$ on $t$, as well as the dependence on $j$ for $P\triangleq P_{j}$, accuracy-confidence function $A(q)\triangleq A_{j}(q)$, and confidence distribution $F(q)\triangleq F_{j}(q)$. Additionally, where appropriate, we drop the dependence of the predictor’s confidence $Q^{\star}(\cdot)$ and of $C(\cdot)$ on $\accentset{\leftarrow}{H}_{j}^{k}{(t)}$ and $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$, respectively. We denote by $Q^{\star}$ a random variable with the same distribution as any confidence level $Q^{\star}$ assigned by the predictor to a resource $j$.
Before moving on to the next section, we reiterate our problem statement here. This paper builds on [1] and substantially expands on it by providing a comprehensive analysis of the calibration behavior of the outage predictor in an ML-assisted resource allocation system. In particular, [1] analyzed the general case where the outage predictor may or may not be well calibrated. In such cases, the predictor’s confidence scores may not accurately reflect the true probability of outages. Central to our study is the characterization of the system’s OP under perfect calibration, which helps guide the choice of the classification threshold required to achieve a target OP when sufficient resources are available. We then apply post-processing calibration techniques to validate our theoretical findings.